# Sensory Data Sharing in ISAC System: Fundamental Limits and Waveform Design

### v2

@@ VII Conclusion
In this paper, we delve into the dual-functional signaling strategy for communication-assisted sensing (CAS) systems across three primary aspects. Firstly, we develop a modified source-channel separation theorem (MSST) tailored for CAS systems under the condition of separable distortion metric. We elucidate the operational meaning of the proposed MSST from the perspective of coding theory. Secondly, we develop an input distribution optimization scheme for the CAS system under scalar channels, which minimizes the CAS distortion while adhering to the MSST and resource constraints. We determine the optimal distribution for CAS systems, balancing between the conventional communication-optimal (Gaussian distribution) and sensing-optimal (2-ary pulse amplitude modulation) schemes. Thirdly, we propose an ISAC waveform design method for Gaussian signaling in MIMO CAS systems, where a successive convex approximation (SCA) algorithm is conceived to solve for the optimal covariance matrix. Simulation results show the unique performance tradeoff between the sensing and communication processes.

### v3

@@ VII Conclusion
In this paper, we delve into the dual-functional signaling strategy for simultaneous sensing data acquisition and sharing (SDAS) systems across three primary aspects. Firstly, we develop a modified source-channel separation theorem tailored (MSST) for SDAS systems under the condition of separable distortion metric. We elucidate the operational meaning of the proposed MSST from the perspective of coding theory. Secondly, we develop an input distribution optimization scheme for SDAS systems under scalar channels, which minimizes the SDAS distortion while adhering to the MSST constraint. We determine the optimal distribution, balancing between the conventional communication-optimal (Gaussian distribution) and sensing-optimal (2-ary pulse amplitude modulation) schemes. Thirdly, we propose an ISAC waveform design method for Gaussian signaling in MIMO SDAS systems, where a successive convex approximation algorithm is conceived to solve for the optimal covariance matrix. Simulation results show the unique performance tradeoff between the sensing and communication processes.